Solve the equation below by using the addition (elimination) method:
3x-y=13
2x+y=3

Found 2 solutions by Alan3354, mananth:
Answer by Alan3354(69443)

3x-y=13
2x+y=3
-------- Add them

Answer by mananth(16949)

3x-y=13
2x+y=3
..
3x+2x-y+y=13+3
5x=16
/5
x=16/5
..
plug value of x in equation 1
3*(16/5)-y=13
48/5 -y =13
y=48/5 -13
y=(48-65)/5
y=-17/5
...
m.ananth@hotmail.ca